# Long-Tail Learning with Rebalanced Contrastive Loss

### v1

@@ 4 Experiments
To validate our findings on representation learning optimization with RCL, we use three benchmark datasets containing significant class imbalance, exhibiting long-tail distribution and involving a large number of classes. We next describe the datasets, experiment design, and performance metrics, followed by comparisons of the results with the SOTA techniques and ablation studies.
@@ 4.1 Datasets
We use the long-tail versions of the benchmark datasets: CIFAR10, CIFAR100, and ImageNet. The level of class imbalance is measured by the Class Imbalance Factor (IF), which is the ratio between the most frequent class frequency and the least frequent class frequency. We set the maximum possible imbalance factors for the datasets, leaving at least 5-6 instances per tail class. Doing so, we obtain the following experiment settings; i) CIFAR10: No. of classes 10, IF - 100, majority class frequency - 5,000, ii) CIFAR100: No. of classes 100, IF - 100, majority class frequency - 500, iii) ImageNet: No. of classes 1,000, IF - 256, majority class frequency - 1,280.
@@ 4.2 Implementation
Referring to the previous implementation, we use ResNet32 backbone for the CIFAR datasets and ResNext-50 and ResNet50 for the ImageNet_Lt. Models are trained with a batch size of 256. Learning rate for CIFAR10 and 100 are set to 0.2 and for ImageNet 0.1. Similar to [7] and [29], we use Autoaugment [6] and Cutout [9] for augmenting the features fed to the classifier. SimAugment [4] is used for the features fed to the contrastive loss. The model training is conducted on the Amazon Cloud with RONIN mediation service with 8xV100 GPUs.
@@ 4.3 Ablation Study
We first conduct an ablation study on CIFAR100 with IF set to 100 to analyse the performance benefits of RCL. The overall balanced accuracy (averaged per class) is reported in Table 1. As small values heavily influence arithmetic means in obtaining the overall class accuracy, we also report the harmonic mean to reflect the performance of the worst-performing classes. Initially, the dataset is tested for cross-entropy loss with LA [20]. Where the logit adjustment parameter is set to 1.3 through grid search. Then, the accuracy is evaluated by combining standard supervised contrastive loss with setting to 2 and to 1 and using SimAugment for augmentation. The compression factor is set to 0.005 for the classes with validation accuracy less than , and for the other classes, the factor is set to 1. Compression is applied from the epoch.
Results in Table 1 show adding RCL to standard SCL would increase top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4). When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5). SCL+RCL demonstrates the highest harmonic mean accuracy of all the combinations.
@@ 4.4.1 Baselines
We compare our proposed framework with twelve SOTA models. Balanced overall test accuracy (arithmetic mean over per-class accuracy) is used for comparison. The selected baselines are: 1) standard cross entropy minimisation, 2) LDAM-DRW where logit margin is enforced depending on the class frequency and reweighting scheme is used in optimization [2], 3) Class balanced focal loss [8] which assigns a frequency based weighting on per class losses, 4) Additive logit margin based on the class frequency [20], 5) BALMS - Balanced SoftMax function Equation [23] to address the distribution shift integrated with a meta sampler that can compute the optimal class sample rate, 6) Decoupled learning which enhances the classification accuracy through fine-tuning a learnt classifier with class-balanced sampling or classifier weight normalization [14], 7) robustness theory based loss to enhance the classification and feature representation [24], 8) RIDE [28] and 9) ACE [1] which are multi expert ensemble models, 10) ELM [12] which enforces logit and embedding margins, and 11) Hybrid_spc [27] prototypical supervised contrastive (PSC) learning.
@@ 4.4.2 Results
We show the results in Table 2. As can be seen, standalone RCL demonstrates , improvement over CIFAR10 and 100, respectively, compared to BCL, which is the next best model. Furthermore, BCL+RCL has consistent improvements , , and over all the datasets in comparison to BCL. In summary, standalone RCL and BCL+RCL outperform all the other baselines apart from one scenario where BCL performs best (i.e., ImageNet).
@@ 4.4.3 Embedding space analysis
In Figure 3, we show the TSNE plots of the learnt representation of the model for CIFAR10 and ImageNet. BCL+RCL demonstrates a better class separation and balanced division of the feature space compared to BCL. As seen from Figure 3(a), in BCL, tail class embeddings are grouped into sub-clusters and lie far apart. Once the features compression is added, all embeddings corresponding to the same tail class are grouped together, providing better intra-class compactness and balancedness in the space in Figure 3(b). For example, in Figure 3(a), tail classes 8 and 9, represented by orange and cyan, have two sub-clusters each, and in Figure 3(b), they have concatenated to two main clusters. Furthermore, when Figures 3(c) and (d) are compared, cluster separability has increased, and all classes occupy the feature space more equally compared to BCL when RCL is applied.
@@ 5 Conclusion
The paper proposed RCL for long-tail classification tasks. RCL addresses the balancedness of the learnt feature space, enforces feature compactness and increases the margin width for tail classes. RCL shows a competitive level of accuracy over SOTA methods over three benchmark datasets with imbalance ratios up to 256. RCL was also validated to perform well as a standalone and integrated loss on BCL, an existing supervised contrastive learning loss specifically designed for long-tail classification. It would be an important future work to investigate how much RCL can contribute to optimizing the learnt feature distribution with self-supervised contrastive learning under extreme class imbalance (i.e., higher class imbalance factors than 256).

### v2

@@ 4 Experimental Setup and Results
To validate our findings on representation learning optimization with RCL, we use three benchmark datasets with significant class imbalance, exhibiting long-tail distribution and involving a large number of classes. Next, we describe the datasets, experiment design followed by comparisons of the results with the SOTA and ablation study.
@@ 4.1 Datasets
We use the long-tail versions of the benchmark datasets: CIFAR10_Lt, CIFAR100_Lt, and ImageNet_Lt. The level of class imbalance is measured by the Class Imbalance Factor (IF), which is the ratio between the most frequent class frequency and the least frequent class frequency. We set the maximum possible imbalance factors for the datasets, leaving at least 5-6 instances per tail class. Doing so, we obtain the following experiment settings; i) CIFAR10_Lt: No. of classes 10, IF - 100, majority class frequency - 5,000, ii) CIFAR100_Lt: No. of classes 100, IF - 100, majority class frequency - 500, iii) ImageNet_Lt: No. of classes 1,000, IF - 256, majority class frequency - 1,280.
@@ 4.2 Implementation
Referring to the previous implementation, we use ResNet32 backbone for the CIFAR datasets and ResNext-50 and ResNet50 for the ImageNet_Lt. Models are trained with a batch size of 256. Learning rate for CIFAR10_Lt and CIFAR100_Lt are set to 0.2 and for ImageNet_Lt 0.1. Similar to work of Cui et al. (2021) and Zhu et al. (2022), we use Autoaugment (Cubuk et al. (2019)) and Cutout (DeVries & Taylor (2017)) for augmenting the features fed to the classifier. SimAugment (Chen et al. (2020)) is used for the features fed to the contrastive loss. The model was trained on a Amazon Cloud VM with 8xV100 GPUs.
@@ 4.3 Ablation Study
We conducted an ablation study on CIFAR100_Lt and CIFAR10_Lt with IF set to 100 to analyse the performance benefits of RCL. The overall balanced accuracy (averaged per class) is reported in Table 1 and Table 2, respectively. As large values heavily influence arithmetic means in obtaining the overall class accuracy, we also report the harmonic mean to reflect the performance of the worst-performing classes. Initially, the dataset is tested for cross-entropy loss with LA (Menon et al. (2020)). Where the logit adjustment parameter is set to 1.3 for RCL through grid search. Then, the accuracy is evaluated by combining standard supervised contrastive loss with setting to 2 and to 1 and using SimAugment for augmentation. The compression factor is set to 1 for the classes with validation accuracy less than ; for the other classes, the factor is set to 0.005. Compression is applied from the epoch.
CIFAR100_Lt results in Table 1 show adding RCL to standard SCL would increase top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4). When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5). SCL+RCL demonstrates the highest harmonic mean accuracy of all the combinations.
Similarly, Table 2 shows the results for CIFAR10_Lt where adding RCL to SCL increases top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4). When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5). SCL+BCL+RCL demonstrates the highest harmonic mean accuracy of all.
@@ 4.4.1 Baselines
We compare our proposed framework with 13 SOTA models under five main categories. Balanced overall test accuracy (arithmetic mean over per-class accuracy) is used for comparison. The selected categories of algorithmic level solutions are 1) Standard and modified (logit adjusted) CE minimization, 2) Two-stage learning methods where representation learning and classifier fine-tuning are handled in different stages; 3) Ensemble methods where multiple experts combined to enhance the head and tail class accuracy simultaneously, 4) CE loss combined with an additional loss to handle representation learning explicitly, 5) CE loss combined with SCL loss to enhance representation learning.
@@ 4.4.2 Results
We show the results in Table 4. As can be seen, standalone RCL demonstrates satisfactory performance compared to the SOTA methods. The improvement over BCL is 1.7%,1.3% and 0.6% for CIFAR10_Lt for IF = 100, 50 and 10, respectively. Similar to CIFAR100_Lt, the values are 0.3%,0.2% and 0.7%. Even though the CIFAR100_Lt improvement over BCL is small compared to the other popular SCL methods, RCL has a competitive performance. Furthermore, BCL+RCL has consistent improvements 1.9%, 1.2%, and 0.9% for CIFAR10_Lt. And for CIFAR100_Lt, the corresponding improvements are 0.2% for IF100 and 0.5 % IF=10. In summary, BCL performs better for CIFAR100_Lt IF=50. For Imagenet_Lt, BCL+RCL outperforms BCL by 0.1% in both backbones according to Table 4.
| Dataset (IF=100) | BCL | BCL+RCL | ||
| CHI | DBI | CHI | DBI | |
| CIFAR100_Lt | 3,962.9 | 0.97 | 4,091.8 | 0.94 |
| CIFAR10_Lt | 84.1 | 4.1 | 105.8 | 4.0 |
@@ 4.4.3 Embedding Space Analysis
According to the quantitative analysis of the RCL method performance, it is evident that standalone RCL applied to the standard SCL demonstrates SOTA competitive accuracy levels. Apart from this, RCL demonstrates an interesting aspect when applied to the BCL. Even though the accuracy improvement over BCL is relatively small, adding the RCL to the BCL enhances the learnt feature distribution. In Figure 3, we show the TSNE plots of the learnt representation of the model for CIFAR10_Lt and CIFAR100_Lt. BCL+RCL demonstrates a better class separation and balanced feature space division than BCL. As seen from Figure 3(a), in BCL, tail class embeddings are grouped into sub-clusters and are far apart. Once the features compression is added, all embeddings corresponding to the same tail class are grouped together, providing better intra-class compactness and balancedness in the space. Figure 3(b) visually showcase the better class separation and balanced space distribution for all classes except for 1 and 9 compared to Figure 3(a); furthermore, when Figures 3(c) and (d) are compared, cluster balancedness has increased as a whole, as all classes occupy the feature space more equally compared to BCL when RCL is applied. In addition, compactness for classes 40 and 90 appears to increase significantly. To further validate this phenomenon, we compute the Calinski–Harabasz index (Caliński & Harabasz (1974)) for the learnt features of the validation set. A higher index value indicates better cluster cohesion and cluster separation. Table 5 indicates the index values for CIFAR10_Lt and CIFAR100_Lt. Where we can see a consistent increment in the Calinski–Harabasz index for both datasets for BCL+RCL over BCL. Davies-Bouldin Index assesses the clustering quality where a lower value indicates more compact and well-separated clusters. According to the table, we can see that both datasets have a consistently lower value for BCL+RCL over BCL, which implies better cluster quality. The index is relatively insensitive to the choice of distance metric. Based on this analysis, we can conclude that RCL has the capability to optimize the distribution of the learnt feature simultaneously with the classification accuracy. The feature distribution is improved in balancedness, class compactness (cohesion) and regularization (larger margins, better class separation).
@@ 5 Conclusion
The paper proposed RCL for long-tail classification tasks. RCL addresses the balancedness of the learnt feature space, enforces feature compactness and increases the margin width for tail classes. RCL shows a competitive level of accuracy over SOTA methods over three benchmark datasets with imbalance factors up to 256. RCL was also validated to perform well as a standalone and integrated loss on BCL, an existing supervised contrastive learning loss specifically designed for long-tail classification. It would be an important future work to investigate how much RCL can contribute to optimizing the learnt feature distribution with self-supervised contrastive learning under extreme class imbalance (i.e., IF than 256).
Acknowledgments
This research was conducted by the University of Sydney for National Intelligence Postdoctoral Grant (project number NIPG-2022-006) and funded by the Australian Government.